# RoboMamba: Multimodal State Space Model for Efficient Robot Reasoning and Manipulation

### v1

@@ 4.1 Experiment settings
Datasets (Stage 1) In the alignment pre-training stage, we utilize the LLaVA-LCS 558K dataset [59], which is a curated subset of the LAION-CC-SBU dataset, supplemented with captions. During the instruction co-training stage, we created a combined dataset totaling 1.8 million samples, including LLaVA-v1.5 655K mix [59], LRV-INSTRUCT 400K [60], and RoboVQA 800K dataset [27]. The detailed descriptions of the datasets is shown in Appendix B.
@@
Datasets (Stage 2) For the dataset used in the robot manipulation fine-tuning stage, we utilize the SAPIEN [28] to set up an interactive simulation environment with articulated objects from PartNet [45]. To generate data, we use the Franka Panda Robot to randomly interact with objects. When successful manipulation occurs, we record the success 6-DOF poses of the end-effector, which serves as ground truth labels for training. In the training set, we collect 10K images across 20 categories. For testing, we use a set of 1.1K images that include both seen categories from the training set and unseen categories. The details of the categories are provided in Appendix B.
@@
Implementation details Before training, RoboMamba loads a pre-trained CLIP/SigLIP ViT-Large [26, 61] as the visual encoder, and the 2.8/1.4B Mamba [1] model as the language model. During the alignment pre-training and instruction co-training, we conduct training for 1 epoch and 2 epochs, respectively. We utilize the AdamW optimizer with and a learning rate (LR) of 2e-5. The precision of floating-point calculations is set to 16-bit. For manipulation fine-tuning, we train the model for 5 epochs, setting the LR to 1e-5 and applying a weight decay of 0.1. The floating-point precision is set to 32-bit. All experiments are conducted on NVIDIA A100 GPUs.
@@
Reasoning evaluation benchmarks To evaluate reasoning capabilities, we employ several popular benchmarks, including VQAv2 [62], OKVQA [63], RoboVQA [27], GQA [64], OCRVQA [65], VizWiz [66], POPE [67], MME [68], MMBench [69], and MM-Vet [70]. As detailed in Appendix E, we describe the key aspects each benchmark focuses on when assessing models in the field of robotics. Notably, we also directly evaluate RoboMamba’s robot-related reasoning abilities on the 18k validation dataset of RoboVQA, covering robotic tasks such as long-horizon planning, success classification, discriminative and generative affordance, past description, and future prediction.
@@
Manipulation evaluation benchmarks To evaluate our model’s manipulation capabilities, we follow previous works [44, 54, 15] and test pulling accuracy exclusively in the simulator [28]. We use the predicted contact point and rotation to interact with objects. To measure the model’s performance, we use the classical manipulation success rate, defined as the ratio of successfully manipulated samples to the total test samples. A manipulation action is considered successful if the difference in the object’s joint state before and after interaction exceeds a threshold of 0.1 meters. In real-world experiments, we use the Franka Panda robot to manipulate several articulated objects.
@@
Method LLM Size Res. OKVQA VQAV2 GQA VizWiz OCR-VQA POPE MME MMB MM-Vet BLIP-2 [49] 7B 224 45.9 - 41.0 19.6 40.6 85.3 1293.8 - 22.4 InstructBLIP [71] 7B 224 - - 49.5 33.4 44.8 - - 36 26.2 LLaMA-AdapterV2 [51] 7B 336 49.6 70.7 45.1 39.8 - - 1328.4 - - MiniGPT-v2 [72] 7B 448 57.8 - 60.1 53.6 - - - - - Qwen-VL [73] 7B 448 58.6 79.5 59.3 35.2 75.7 - - 38.2 - LLaVA1.5 [59] 7B 336 - 78.5 62.0 50.0 - 85.9 1510.7 64.3 30.5 SPHINX [55] 7B 224 62.1 78.1 62.6 39.9 66.0 80.7 1476.1 66.9 36.0 LLaVA-Phi [74] 2.7B 336 - 71.4 - 35.9 - 85.0 1335.1 59.8 28.9 MobileVLM [75] 2.7B 336 - - 59.0 - - 84.9 1288.9 59.6 - TinyLLaVA [76] 2.7B 336 - 77.7 61.0 - - 86.3 1437.3 68.3 31.7 RoboMamba(Ours) 2.7B 224 63.1 80.3 62.4 55.0 62.5 85.3 1314.8 64.2 28.6 RoboMamba(Ours) 2.7B 384 62.4 79.1 64.4 55.0 66.7 86.9 1354.2 65.7 29.7
4.2 Reasoning quantitative result
@@
General reasoning. As shown in Table 1, we compare RoboMamba with previous state-of-the-art (SOTA) MLLMs on general VQA and recent MLLM benchmarks. First, we find that RoboMamba achieves promising results across all VQA benchmarks, using only a 2.7B language model. The results demonstrate that our simple architecture design is effective. The alignment pre-training and proposed instruction co-training significantly enhance the MLLM’s reasoning capabilities. For example, due to the large amount of robot data introduced during the co-training stage, our model’s spatial identification performance on the GQA benchmark is improved. Meanwhile, we also test our RoboMamba on recently proposed MLLM benchmarks. Compared to previous MLLMs, we observe that our model achieves competitive results across all benchmarks. Notably, our model achieves satisfactory results on the POPE benchmark due to the inclusion of the LRV-Instruct dataset during the co-training stage, which helps reduce failed robot actions caused by hallucinations. Although some performances of RoboMamba are still below those of LLaVA1.5 and SPHINX, we prioritize using a smaller and faster Mamba to balance the efficiency of the robotic model. In the future, we plan to develop RoboMamba-7B for scenarios where resources are not limited.
@@
Robot-related reasoning. To comprehensively compare RoboMamba’s robot-related reasoning abilities, we benchmark it against LLaMA-AdapterV2 [51] on the RoboVQA [27] validation set. We chose LLaMA-AdapterV2 as a baseline because it serves as the base model for the current SOTA Robot MLLM, ManipLLM [15]. For a fair comparison, we loaded the baseline pre-trained parameters and fine-tuned it on the RoboVQA training set for two epochs, using its official instruction tuning method. As shown in Figure 3 a), RoboMamba achieves superior performance across BLEU-1 to BLEU-4. The results indicate that our model possesses advanced robot-related reasoning capabilities and confirms the effectiveness of our training strategy. In addition to higher accuracy, our model achieves inference speeds 7 times faster than LLaMA-AdapterV2 and ManipLLM, which can be attributed to the content-aware reasoning ability and efficiency of the Mamba language model [25]. Finally, we visualize the qualitative results in Figure 4.
@@
4.3 Manipulation quantitative result
Results. As shown in Table 2, our RoboMamba achieves a 7.0% improvement on seen categories and a 2.0% improvement on unseen categories compared to the previous SOTA ManipLLM. Moreover, our method showcases SOTA performance across 14 of 20 seen categories, highlighting its effectiveness and stability in predicting action poses. For unseen categories, the recent three MLLM-based methods—RoboFlamingo, ManipLLM, and our method—all achieved promising performance. The results demonstrate that leveraging the strong generalization abilities of MLLMs can effectively improve the policy’s generalization ability while enhancing accuracy on unseen objects. Regarding efficiency, RoboFlamingo updates 35.5% (1.8B) of the model parameters, ManipLLM updates an adapter (41.3M) comprising 0.5% of the model parameters, whereas our fine-tuned simple policy head (3.7M) only constitutes 0.1% of the model parameters. RoboMamba effectively updates 10 times fewer parameters than previous MLLM-based methods while achieving seven times faster inference speeds. The results reveal that our RoboMamba not only possesses strong reasoning abilities but also can acquires manipulation capabilities in a cost-effective manner.
@@ 4.4 Ablation study
The impact of reasoning ability. We explore whether utilizing MLLMs with different reasoning abilities affects manipulation skill learning. For a fair comparison, we use the same manipulation fine-tuning strategy, injecting and fine-tuning a simple MLP policy head after the MLLM (while freezing other parameters). We compare our RoboMamba-2.7B (Ours-2.7B) with OpenFlamingo, LLaMA-AdapterV2, and our RoboMamba-1.4B. As shown in Figure 3 b), Ours-2.7B achieves promising results compared with other methods, which is proportional to its reasoning ability. Meanwhile, Ours-2.7B (w/o C) indicates that we did not use the instruction co-training method, omitting the robot-related RoboVQA dataset during fine-tuning. We find that this also impacts the accuracy of manipulation, especially reducing the model’s generalization ability when facing unseen objects. The results confirm our finding: fine-tuning an MLLM to learn robot skills does not require extensive resources; it only requires that the MLLM possesses strong robot-related reasoning abilities. Additionally, we present more ablation studies in Appendix C, including explorations of training strategies and policy head design.
@@ 4.5 Real-world experiments
As shown in Figure 4, we visualize RoboMamba’s reasoning results across various robotic downstream tasks. For task planning, compared to LLaMA-AdapterV2, RoboMamba demonstrates more accurate and long-horizon planning abilities, thanks to its strong reasoning capabilities. For a fair comparison, we also fine-tuned the baseline LLaMA-AdapterV2 on the RoboVQA dataset. Additionally, RoboMamba accurately performs fundamental robotic tasks such as affordance generation and discrimination, proving that it can understand robotic scenes. Notably, our model also possesses past and future prediction capabilities, further highlighting its robust reasoning capabilities. For pose prediction, we use a Franka Emika robotic arm to interact with various household objects. We project RoboMamba’s predicted 3D pose back onto a 2D image, using a red dot to indicate the contact point and the end-effector to show the direction, as shown in the bottom right corner of the figure. More real-world demonstrations are provided in Appendix 5 and the supplementary video file.
@@ 5 Conclusion and future plan
In this paper, we introduce an end-to-end robotic MLLM named RoboMamba, which possesses both reasoning and manipulation capabilities. Based on our RoboMamba, we can impart new manipulation skills to the model by fine-tuning a simple policy head (0.1% of the model) in approximately 20 minutes. This finding reveals how to efficiently equip an MLLM with manipulation abilities without compromising its inherent reasoning capabilities. Finally, RoboMamba excels in reasoning on both general and robot-related evaluation benchmarks and showcases impressive pose prediction results. As for our future plan, we focus on two directions. 1) During the process for robot manipulation fine-tuning, we aim to introduce continual learning techniques [77, 78, 79] to continuously enhance the model’s manipulation abilities in the real world. 2) Constructing a 3D Robot MLLM [80, 81, 82], as the 3D point cloud contains more geometric information that helps predict 3D manipulation poses.
@@ Appendix A Appendix
Due to space limitations, we provide additional details of the proposed method in this supplementary material. In Appendix B, we offer a more detailed description of our training dataset, including alignment pre-training, instruction co-training, and robot manipulation fine-tuning. Additional ablation study are presented in Appendix C, which explore the impact of different training strategies on reasoning ability and the effect of different head designs on manipulation fine-tuning. In Appendix 5, we show additional qualitative results across multiple robot-related downstream tasks. Finally, we provide the metric selection rationale and the usage of prompts during testing in Appendix E.
@@ Appendix B Dataset description
2) LRV-INSTRUCT 400K: Although recent work [55] suggests that increasing image resolution can effectively mitigate hallucination, processing high-resolution images affects the efficiency of robotic policy. Therefore, we introduced this dataset during the co-training process to mitigate hallucination. This dataset contains visual instructions generated by GPT-4, encompassing 16 different vision-and-language tasks with open-ended instructions and answers.
3) RoboVQA 800K: In co-training, we use this dataset to enhance our model’s robot-related reasoning abilities. RoboVQA [27] comprises realistic data collected by performing various user requests and using multiple embodiments, such as robots, humans, and humans with grasping tools. This dataset includes 5,246 long-horizon episodes and 92,948 medium-horizon episodes of robotic tasks, each paired with image and text prompt inputs.
@@
Representation for Each Category Icon In Table 3, we provide an overview of the meaning of each category icon presented in Table 2 of the main paper. These categories and their corresponding objects are sourced from PartNet-Mobility [89].
Simulator Data Collection In the simulator, we use a Franka Panda Robot with a suction gripper as the robotic actuator. During data collection, we randomly select a contact point on the movable part of the object and orient the end-effector’s z-axis opposite to the object’s normal vector, with a random y-axis direction to interact with the object. Successful operations are categorized as successful samples and integrated into the training dataset. For the training set, we collect 10K images across 20 categories, including Safe, Door, Display, Refrigerator, Laptop, Lighter, Microwave, Mouse, Box, Trash Can, Kitchen Pot, Suitcase, Pliers, Storage Furniture, Remote, Bottle, Folding Chair, Toaster, Lamp, and Dispenser. For testing, we use a set of 1.1K images that include both seen categories from training and unseen categories, such as Toilet, Scissors, Table, Stapler, Kettle, USB, Switch, Washing Machine, Faucet, and Phone.

### v2

@@ 4.1 Experiment settings
Datasets (Stage 1) In the alignment pre-training stage, we utilize the LLaVA-LCS 558K dataset [67], which is a curated subset of the LAION-CC-SBU dataset, supplemented with captions. During the instruction co-training stage, we combine general instruction datasets with the robotic instruction datasets. Specifically, for the general instruction dataset, we selectively adopt the LLaVA mixed instruction dataset [4], the ShareGPT4V-SFT dataset [68], or the LLaVA-Next dataset [69]. For the robotic instruction dataset, we randomly sample some image-text paired training samples from the RoboVQA [27] dataset. In our main experiments, a mixture of the LLaVA 1.5 instruction dataset and the 300K RoboVQA dataset is used during the co-training stage. Detailed descriptions of these datasets are provided in Appendix B.
@@
Datasets (Stage 2) For the dataset used in the robot manipulation fine-tuning stage, we follow the data collection process of previous works [61, 15], adopting the SAPIEN engine [28] to set up an interactive simulation environment with articulated objects from PartNet-Mobility [58]. The Franka Panda Robot, equipped with a suction gripper, serves as the robotic actuator. During data collection, we randomly select a contact point p on the movable part and orient the end-effector’s z-axis opposite to its normal vector, with a random y-axis direction to interact with the object. Successful operations are categorized as successful samples and integrated into the dataset. In the training set, we collect 10K images across 20 tasks. For evaluation, we generate 1.1K examples for the test set, comprising 20 training (seen) and 10 testing (unseen) tasks. The unseen tasks are used to evaluate the generalization capability of our model. The details of the categories are provided in Appendix B.
@@
Implementation details Before training, RoboMamba loads a pre-trained CLIP/SigLIP ViT-Large [26, 70] as the visual encoder, and the 2.8/1.4B Mamba [1] model as the language model. During the alignment pre-training and instruction co-training, we conduct training for 1 epoch and 2 epochs, respectively. We utilize the AdamW optimizer with and a learning rate (LR) of 4e-5. The precision of floating-point calculations is set to 16-bit. For manipulation fine-tuning, we train the model for 8 epochs, setting the LR to 1e-5 and applying a weight decay of 0.1. The floating-point precision is set to 32-bit. All experiments are conducted on NVIDIA A100 GPUs.
@@
Reasoning evaluation benchmarks To evaluate reasoning capabilities, we employ several popular benchmarks, including VQAv2 [71], OKVQA [72], RoboVQA [27], GQA [73], VizWiz [74], POPE [75], MME [76], MMBench [77], and MM-Vet [78]. As detailed in Appendix E, we describe the key aspects each benchmark focuses on when assessing models in the field of robotics. Notably, we also directly evaluate RoboMamba’s robotic-related reasoning abilities on the 18k validation dataset of RoboVQA, covering robotic tasks such as long-horizon planning, success classification, discriminative and generative affordance, past description, and future prediction.
@@
Manipulation evaluation benchmarks To evaluate our model’s manipulation capabilities, we follow previous works [57, 63, 15] and test open-loop task completion accuracy exclusively in the simulator [28]. The predicted contact point and rotation are used to interact with objects. To measure the model’s performance, we use the classical manipulation success rate, defined as the ratio of successfully manipulated samples to the total test samples. A manipulation action is considered successful if the difference in the object’s joint state before and after interaction exceeds a threshold of 0.1 meters. In real-world experiments, we use the Franka Panda robot to manipulate several articulated objects.
@@
Method LLM Res. OKVQA VQAV2 GQA VizWiz POPE MME MMB MM-Vet RoboVQA4 RoboVQA1 BLIP-2 [43] 7B 224 45.9 - 41.0 19.6 85.3 1293.8 - 22.4 - - InstructBLIP [79] 7B 224 - - 49.5 33.4 - - 36 26.2 - - LLaMA-AdapterV2 [45] 7B 336 49.6 70.7 45.1 39.8 - 1328.4 - - 8.1 27.8 MiniGPT-v2 [80] 7B 448 57.8 - 60.1 53.6 - - - - - - Qwen-VL [81] 7B 448 58.6 79.5 59.3 35.2 - - 38.2 - - - LLaVA1.5 [67] 7B 336 - 78.5 62.0 50.0 85.9 1510.7 64.3 30.5 - - SPHINX [64] 7B 224 62.1 78.1 62.6 39.9 80.7 1476.1 66.9 36.0 - - LLaVA-Phi [49] 2.7B 336 - 71.4 35.9 - 85.0 1335.1 59.8 28.9 - - MobileVLM [82] 2.7B 336 - - 59.0 - 84.9 1288.9 59.6 - - - TinyLLaVA [83] 2.7B 336 - 77.7 61.0 - 86.3 1437.3 68.3 31.7 29.6 43.5 RoboMamba(Ours) 2.7B 224 63.3 79.6 64.2 57.1 86.3 1297.2 60.9 29.4 42.8 62.7 RoboMamba(Ours) 2.7B 336 62.7 77.7 63.3 58.1 87.0 1335.5 60.7 31.4 41.8 61.9
4.2 Reasoning quantitative results
@@
General reasoning. As shown in Table 1, we compare RoboMamba with previous state-of-the-art (SOTA) MLLMs on general VQA and recent MLLM benchmarks. First, we find that RoboMamba achieves promising results across all VQA benchmarks, using only a 2.7B language model. The results demonstrate that our simple architecture design is effective. The proposed instruction co-training significantly enhance the MLLM’s reasoning capabilities. For example, due to the large amount of robot data introduced during the co-training stage, our model’s spatial identification performance on the GQA benchmark is improved. Meanwhile, we also test our RoboMamba on recently proposed MLLM benchmarks. Compared to previous MLLMs, we observe that our model achieves competitive results across all benchmarks. Specifically, our model achieves satisfactory results on the POPE benchmark, helping to reduce failed robot actions caused by hallucinations. Although some performances of RoboMamba are still below those of LLaVA1.5 and SPHINX, we prioritize using a smaller and faster Mamba to balance the efficiency of the robotic model. In the future, we plan to develop RoboMamba-7B for scenarios where resources are not limited.
@@
Robotic-related reasoning. To comprehensively compare RoboMamba’s robotic-related reasoning abilities, we benchmark it against LLaMA-AdapterV2 [45] and TinyLLaVA [83] on the RoboVQA [27] validation set. We choose LLaMA-AdapterV2 as a baseline because it serves as the base model for the current state-of-the-art (SOTA) robot MLLM, ManipLLM [15]. Meanwhile, TinyLLaVA is chosen as a representative tiny MLLM, enabling a comparison of robotic-related reasoning abilities. For a fair comparison, we load the pre-trained parameters of both LLaMA-AdapterV2 and TinyLLaVA and fine-tuned the baseline models on the RoboVQA training set for two epochs, using their official instruction-tuning method. As shown in Table 1, RoboMamba achieves superior performance across BLEU-1 to BLEU-4. The results indicate that our model possesses advanced robotic-related reasoning capabilities and confirms the effectiveness of our training strategy. In addition to higher accuracy, our model achieves inference speeds 7 times faster than LLaMA-AdapterV2 and ManipLLM, which can be attributed to the content-aware reasoning ability and efficiency of the Mamba language model [25]. Finally, we visualize the qualitative results in Figure 4.
4.3 Manipulation quantitative results
@@
Results. As shown in Table 2, our RoboMamba achieves a 7.0% improvement on seen tasks and a 2.0% improvement on unseen tasks compared to the previous SOTA ManipLLM. Moreover, our method showcases SOTA performance across 14 of 20 seen tasks, highlighting its effectiveness and stability in predicting action poses. For unseen tasks, the recent three MLLM-based methods—RoboFlamingo, ManipLLM, and our method—all achieved promising performance. The results demonstrate that leveraging the strong generalization abilities of MLLMs can effectively improve the policy’s generalization ability while enhancing accuracy on unseen objects. Regarding efficiency, RoboFlamingo updates 35.5% (1.8B) of the model parameters, ManipLLM updates an adapter (41.3M) comprising 0.5% of the model parameters, whereas our fine-tuned simple policy head (3.7M) only constitutes 0.1% of the model parameters. RoboMamba effectively updates 10 times fewer parameters than previous MLLM-based methods while achieving seven times faster inference speeds. The results reveal that our RoboMamba not only possesses strong reasoning abilities but also can acquires manipulation capabilities in a cost-effective manner.
@@ 4.4 Ablation study
The impact of LLM on reasoning abilities. As shown in Figure 3 a), we explore the impact of different LLMs on general and robotic-related reasoning abilities. Given that efficiency is crucial in robotic tasks and directly affects the practicality of policy models, we compare Mamba-2.7B with other linear complexity LLMs. For all experiments, we utilize the same training data and strategy. Compared with RWKV-3B [24], Mamba-2.7B achieves significant improvements on both common sense and robotic-related reasoning benchmarks. The results demonstrate that the Mamba-2.7B model not only possesses linear complexity but also efficiently acquires strong reasoning abilities through our proposed training strategy. Meanwhile, our proposed RoboMamba VLA framework and training strategy can also be adapted to other, more advanced linear-complexity LLM models.
The impact of reasoning abilities on manipulation accuracy. We explore whether utilizing MLLMs with different reasoning abilities affects manipulation skill learning. For a fair comparison, we use the same manipulation fine-tuning strategy, injecting and fine-tuning a simple MLP policy head after the MLLM (while freezing other parameters). We compare our RoboMamba-2.7B (Ours-2.7B) with OpenFlamingo, LLaMA-AdapterV2, and our RoboMamba-1.4B. As shown in Figure 3 b), Ours-2.7B achieves promising results compared with other methods, which is proportional to its reasoning ability. Meanwhile, Ours-2.7B (w/o C) indicates that we did not use the instruction co-training method, omitting the robotic-related RoboVQA dataset during fine-tuning. We find that this also impacts the accuracy of manipulation, especially reducing the model’s generalization ability when facing unseen objects. The results confirm our finding: fine-tuning an MLLM to learn robot skills does not require extensive resources; it only requires that the MLLM possesses strong robotic-related reasoning abilities. Additionally, we present more ablation studies in Appendix C, including explorations of different vision encoders, training datasets, and policy head design.
@@ 4.5 Real-world experiments
As shown in Figure 4, we visualize RoboMamba’s reasoning results across various robotic downstream tasks. For task planning, compared to LLaMA-AdapterV2, RoboMamba demonstrates more accurate and long-horizon planning abilities, thanks to its strong reasoning capabilities. For a fair comparison, we also fine-tuned the baseline LLaMA-AdapterV2 on the RoboVQA dataset. Additionally, RoboMamba accurately performs fundamental robotic tasks such as affordance generation and discrimination, proving that it can understand robotic scenes. Notably, our model also possesses past and future prediction capabilities, further highlighting its robust reasoning capabilities. Prediction of past and future actions is crucial in robotic manipulation, as it not only enables effective rethinking of past failure actions but also enhances the robustness of future manipulation pose generation. For low-level action, we use a Franka Emika robotic arm to interact with various household objects. Due to the direct visualization of the gripper causing occlusion, we project RoboMamba’s predicted 6 DoF pose onto a 2D image, using a red dot to indicate the contact point and the end-effector to show the rotation, as shown in the bottom right corner of the figure.More real-world demonstrations are provided in Appendix 6 and the supplementary video file. Meanwhile, as shown in Figure 5, we also visualize the failure cases of RoboMamba’s predictions in both reasoning and manipulation tasks.
@@ 5 Conclusion and future plan
In this paper, we introduce RoboMamba, an efficient VLA model that combines a vision encoder with the linear-complexity Mamba LLM, equipped with visual common sense reasoning and robotic reasoning abilities. Based on our RoboMamba, we can impart new manipulation skills to the model by fine-tuning a simple policy head (0.1% of the model) in a few dozen minutes. This finding reveals how to efficiently equip an VLA model with manipulation abilities without compromising its inherent reasoning capabilities. Finally, RoboMamba excels in reasoning on both general and robotic-related evaluation benchmarks and showcases impressive pose prediction results. Regarding limitations, while our proposed RoboMamba achieves efficient inference speed, its reliance on a 2.7B LLM leads to limitations on certain complex reasoning tasks when compared to MLLMs built on 7B/13B LLMs. Looking ahead, our future work will focus on two main directions. 1) We plan to adapt the RoboMamba VLA framework and training strategy to more advanced linear-complexity LLM models to further enhance its reasoning and manipulation capabilities. 2) Constructing a 4D Robot VLA model [84, 48, 85], as 3D point cloud and temporal data contain more robotics-specific information that aids in predicting robust low-level actions.
Acknowledgements. This work was supported by the National Natural Science Foundation of China (62476011).
@@ Appendix A Appendix
Due to space limitations, we provide additional details of the proposed method in this supplementary material. In Appendix B, we offer a more detailed description of our training dataset, including alignment pre-training, instruction co-training, and robot manipulation fine-tuning datasets. Additional ablation studies are presented in Appendix C, exploring the impact of different image encoders on reasoning ability, the impact of different training datasets on reasoning ability, and the effect of various head designs on manipulation fine-tuning. In Appendix 6, we show additional qualitative results across multiple robot-related downstream tasks. Finally, we provide the metric selection rationale and the usage of prompts during testing in Appendix E.
@@ Appendix B Dataset description
2) ShareGPT4V-SFT dataset: The ShareGPT4V-SFT dataset is similar to LLaVA-v1.5 655K [67], except that the 23K detailed description entries in LLaVA-1.5-SFT are replaced with detailed captions randomly sampled from the 100K ShareGPT4V data [86].
3) LLaVA-Next dataset: Compared to LLaVA-v1.5 655K [67], LLaVA-Next [69] enhances the instruction data mixture by prioritizing high-quality user instruction data and expanding multimodal document/chart data sources. For high-quality user instruction data, LLaVA-Next ensure task diversity and response quality by using existing GPT-V data (LAION-GPT-V and ShareGPT-4V) and a carefully curated 15K visual instruction dataset from LLaVA demos. Additionally, LLaVA-Next replaces TextCaps with DocVQA and SynDog-EN to improve OCR capability.
4) RoboVQA 800K: In co-training, we use this dataset to enhance our model’s robot-related reasoning abilities. RoboVQA [27] comprises realistic data collected by performing various user requests and using multiple embodiments, such as robots, humans, and humans with grasping tools. This dataset includes 5,246 long-horizon episodes and 92,948 medium-horizon episodes of robotic tasks, each paired with image and text prompt inputs. In our experiments, we randomly select 300K image-text paired instruction samples from RoboVQA to construct the co-training dataset.
@@
Representation for Each Category Icon In Table 3, we provide an overview of the meaning of each category icon presented in Table 2 of the main paper. Following Partnet [58], different tasks are designed for each category. For instance, opening the door or control panel of a refrigerator, opening the cap of a bottle, and rotating the lid of a box. The detailed task design can be found at: https://sapien.ucsd.edu/browse.
Simulator Data Collection In the simulator, we use a Franka Panda Robot with a suction gripper as the robotic actuator. During data collection, we randomly select a contact point on the movable part of the object and orient the end-effector’s z-axis opposite to the object’s normal vector, with a random y-axis direction to interact with the object. Successful operations are categorized as successful samples and integrated into the training dataset. For the training set, we collect 10K images across 20 categories, including Safe, Door, Display, Refrigerator, Laptop, Lighter, Microwave, Mouse, Box, Trash Can, Kitchen Pot, Suitcase, Pliers, Storage Furniture, Remote, Bottle, Folding Chair, Toaster, Lamp, and Dispenser. For testing, we use a set of 1.1K images that include both seen categories from training and unseen categories, such as Toilet, Scissors, Table, Stapler, Kettle, USB, Switch, Washing Machine, Faucet, and Phone. Regarding the variation between training and testing data, we followed the data collection settings of where2act [61] and ManipLLM [15]. The specific variations can be divided into two aspects: 1) Asset Variation and 2) State Variation.
1) Asset Variation: We use 20 categories from PartNet [58] for seen objects and reserve the remaining 10 categories for unseen objects to analyze if RoboMamba can generalize to novel categories. Specifically, we further divide the seen objects into 1037 training shapes and 489 testing shapes, using only the training shapes to construct the training data. Thus, the shapes of the seen objects encountered during training and testing are different. For unseen categories, there are a total of 274 shapes, which are used exclusively in the testing data.
2) State Variation: We observe the object in the scene from an RGB-D camera with known intrinsics, mounted 4.5-5.5 units away from the object, facing its center. The camera is located at the upper hemisphere of the object with a random azimuth between [-45, 45] and a random altitude between [30, 60]. Since the tasks involve ’pulling,’ we also initialize the starting pose for each articulated part randomly between its rest joint state (fully closed) and any position up to half of its joint state (half-opened). These state settings are utilized for both training and testing data, aiming to boost the model’s generalization ability.